Grade 9Magical Mathematics[Interesting Approach]

solve root(x+16) - root x = 6/root x pls show evrey step!!!!!!!!

Shivam Bhagat
15 Years agoGrade 9

2 Answers

vikas askiitian expert
15 Years ago

(x+16)1/2  - x1/2  = 6/x1/2 

multiply this equation by x1/2

[x(x+16)]1/2 - x = 6

 [x(x+16)]1/2  = x+6

squaring both sides

 x(x+16) = x2 + 12x + 36

 x2 + 16x = x2 +12x +36

      4x = 36

       x = 9

this is the required solution

charitha .bvs
15 Years ago

ROOT(X+16)=6/ROOTX+ROOTX

SQUARE ON BOTH SIDES

X+16=36/X+X+12

36/X=4

X=9